According to the 1998 Guinness Book of World Records stuntman Dan Koko fell a distance of 355 feet into an airbag after jumping from the Vegas World Hotel and Casino. The distance d in feet traveled by a freefalling object in t seconds is given by the formula d = 16t 2. To the nearest tenth of a second, how long did the stuntman's freefall last?

Answer by stanbon(75887)

16t^2 = 355
t^2 = 355/16
t = (1/4)sqrt(355)
==============
Cheers,
Stan H.